OpenStudy (anonymous):

How do I sole ln(x^(2) - x - 2) = 2 + ln(x+1) This is as far as I have gotten ln((x^(2) - x - 2)/(x+1)) = 2 (x^(2) - x - 2)/(x+1) = e^(2) then I divided because I was too lazy to do algebra (x+1)(x-2) = e^(2) = x^(2)-x-2-e^(2) = 0 I'm not allowed to use a calculator I was thinking I could just sub these numbers into the quadratic formula to come up with solutions would that be the right course of action?

OpenStudy (anonymous):

\[x^2-x-2=(x+1)(x-2)\] so you can cancel either first or last

OpenStudy (anonymous):

The script displaying Latex type is crashing blah

OpenStudy (anonymous):

probably easier to do it first, \[\ln((x+1)(x-2))=\ln(x+1)+\ln(x+2)=2+\ln(x+2)+\] \[\ln(x-1)=2\] \[x-2=e^2\] etc

OpenStudy (anonymous):

if you can't see the latex i say it in english the x + 1 cancels, you don't get a quadratic

OpenStudy (anonymous):

I will brb going to restart my browser so I can read what you wrote

OpenStudy (anonymous):